Prove that every finite ring is noetherian.


1
Expert's answer
2021-06-04T05:10:43-0400

Solution:

The statement on localizations follows from the fact that any ideal "J \subset S^{-1} R" is of the form "I \cdot S^{-1} R" . Any quotient "R \/ I" of a Noetherian ring R is Noetherian because any ideal "\bar{J} \subset R / I" is of the form "J \/ I" for some ideal "I \subset J \subset R" . Thus it suffices to show that if R is Noetherian so is R[X] . Suppose "J_{1} \subset J_{2} \subset \ldots" is an ascending chain of ideals in R[X]. Consider the ideals "I_{i, d}" defined as the ideal of elements of R which occur as leading coefficients of degree d polynomials in "J_{i}" . Clearly "I_{i, d} \subset I_{i^{\prime}, d^{\prime}}" whenever "i \leq i^{\prime}" and "d \leq d^{\prime}" . By the ascending chain condition in R there are at most finitely many distinct ideals among all of the "I_{i, d}" . (Hint: Any infinite set of elements of "\mathbf{N} \times \mathbf{N}" contains an increasing infinite sequence.) Take "i_{0}" so large that "I_{i, d}=I_{i_{0}, d} \forall i \geq i_{0}" and all d. Suppose "f \in J_{i}" for some "i \geq i_{0}" . By induction on the degree "d=\operatorname{deg}(f)" we show that "f \in J_{i_{0}}" . Namely, there exists a "g \in J_{i_{0}}" whose degree is d and which has the same leading coefficient as f. By induction "f-g \in J_{i_{0}}" and we proved.